QUESTION Mr JABARI is the Director-General of WBG Ltd an investment private company with 100 shareholders. Last week Mr JabarInvesting in electronics production business which is expected to have huge sales due to restrictions imposed by the GovernmeRegaina Calarlate Price & Share. ii) Show othe Relevant that Can be Calarlated
Answer #1

Year Cash Flow PV of Cash Flo Discounting F w actor (cash flow*dis [1/(1.15^yea counting facto r) 2 3 80+10%= 88+10%= 96.8+10

(ii)

Year 3 Year 4 Year 5 144000000 | 172800000 207360000 57600000 69120000 82944000 34560000 41472000 49766400 Year 1 Year 2 Sale

(i) As Expected Share Price is 995.59 i.e. < Current Actual Price i.e. 1000, Purchasing share is NOT FEASIBLE

(ii) As NPV is Positive, Investing in the project IS FEASIBLE

(If this was helpful then please rate positively. Thank You:)